OpenStudy (anonymous):

The legs of a right triangle are 4 units and 7 units long. To the nearest whole number, what is the length of the hypotenuse? a) 8 units b) 9 units c) 10 units d) 11 units

OpenStudy (anonymous):

Do you remember the Pythagorean Theorem?

OpenStudy (anonymous):

\[a^2+b^2=c^2\]

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

The formula I put there is the Pythagorean Theorem. You fit in the sides for a and b so \[4^2+7^2=c^2\]

OpenStudy (anonymous):

now square the two numbers

OpenStudy (anonymous):

16 + 39?

OpenStudy (anonymous):

7*7=49, but close. Now the equation is \[16+49=c^2\]

OpenStudy (anonymous):

add the two numbers together

OpenStudy (anonymous):

then I square root it ?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so my answer is 8 units?

OpenStudy (anonymous):

rounded up, yes.

OpenStudy (anonymous):

omg, thank you <3

OpenStudy (anonymous):

your welcome, good job! :D